How many years does Steven need to invest his $3,000 at 7% to earn $210 in simple interest?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

How many years does Steven need to invest his $3,000 at 7% to earn $210 in simple interest?
------
Interest = Prt
210 = 3000*0.07t
t = 210/(210)
time = 1 year
========================
Cheers,
Stan H.

Answer by richwmiller(17219)

i=prt
210=3000*.07*t
solve for t
210=3000*7/100*t
210=30*7*t
210=210*t
1=t=one year
check
7/100*3000*1=210
7*30=210
210=210